OpenStudy (anonymous):

Find the area of the regular polygon. Give the answer to the nearest tenth. pentagon with radius 8 m A) 304.3 m2 B) 152.2 m2 C) 30.4 m2 D) 154.2 m2

OpenStudy (asadkarim7):

use the formula (x-2)x180 to find the sum of interior angles which wold be equal to 540 deg divide the pentagon into five triangles with interior angle 108 degrees apply 1/2 (8)(8) sin 108 to find the area of 1 triangle =30.43 m ^2 multiply this answer with 5 to get the area of pentagon 152.17 m^2 therefore 152.2 m^2 answer is B

OpenStudy (anonymous):

why would it need to be equal to 540 degrees?

OpenStudy (anonymous):

okay it doesn't need to be it just is gotcha

OpenStudy (anonymous):

The area of a regular polygon of n sides inscribed in a circle of radius r is \[\text{Area}= n r^2 \text{Cos}\left[\frac{\pi }{n}\right] \text{Sin}\left[\frac{\pi }{n}\right]\]=\[80 \sqrt{\frac{5}{8}-\frac{\sqrt{5}}{8}} \left(1+\sqrt{5}\right)=152.169 \]